# Relationship Between Major Stellar Physical Parameters and Normal Mode Frequencies in Accreting White Dwarf Stars

### v1

@@ 2.5 Choice of Modes
The associated pulsations in accreting WDs are generally attributed to be $g$-modes similar to those commonly applied to the isolated WDs. These $g$-modes are excited thermally by the convection zone related to hydrogen and helium ionization zones, as in DAV and DBV stars (Brickhill, 1991; Arras et al., 2006; Saio, 2013). However, some recent studies in the literature suggest that these pulsations could be Rossby modes (Saio, 2019). In our companion paper, we conducted a simultaneous analysis of both $g$ and $r$-modes, demonstrating that within the currently established framework of the convective driving mechanism, the pulsations in accreting WDs should be expected to have a $g$-mode origin (Kumar & Townsley, 2025). We explicitly computed the visibility amplitudes of both $g$ and $r$-modes and highlighted that while a few retrograde $r$-modes exhibit larger visibility, the low-order $g$ modes possess higher frequencies in the star’s frame, thus making them more likely to be driven in the star’s frame and observable.
Following the findings of Kumar & Townsley (2025), with the investigation of the surface amplitudes of $g$ modes in accreting WDs, we focus exclusively on $g$ modes up to the second-lowest order ($\ell=2$) and disregard the $r$ mode significance in such systems in this present analysis. We use GYRE to compute $g$ mode frequencies just as described in Kumar & Townsley (2025).
@@ 3 Results
We now present our seismological calculations for the accreting WD models described in section 2. We consider the $g$ modes in the rapid rotation limit using the observed spin period of GW Lib, 209 s, as a representative value for CV systems (Szkody et al., 2012). This period is much shorter than the hours or days observed in isolated WDs. We conduct analysis within the Traditional Approximation of Rotation (TAR) and Cowling Approximation. We present $g$ mode frequency variation with $T_{\rm c}$ in section 3.1, quantifying the mode frequencies with and without a partial solid core. Furthermore, we propose a potential new method to identify the mode orders based on the time variation of frequencies during the months of cooling after the strong accretion event. Section 3.2 introduces the variation of the $g$ mode frequencies with the accreted layer of the star, $M_{\rm acc}$. Finally, we show the changes for the different mass models.
@@
Here we begin by showing the $g$ mode eigenfrequencies following the strong accretion event. These strong accretion outbursts occur on relatively short timescales compared to the overall age of the star. However, the intermittent accretion events with finite quiescence periods can significantly disrupt the outer layers, leading to modifications in the $g$ mode frequencies. We investigate the radial mode orders for the dipole mode ($\ell=1$) after the strong accretion event, tracking the evolution of mode frequencies over a period of just over four years. These calculations build upon the work of Kumar & Townsley (2023), but with a significantly improved WD model as outlined in the section 2.2. Figure 4 illustrates the time variations of the $g$ mode frequencies for radial orders $n_{\rm g}=2-16~(top~to~bottom)$ in the observer’s frame of a dipole mode, following the accretion outbursts for a 0.78 $M_{\odot}$ WD model. The left panel corresponds to $T_{\rm c}=5\times 10^{6}$ K, where the WD core has a solid core, while the right panel shows results for $T_{\rm c}=7\times 10^{6}$ K. The dashed blue lines are the prograde modes ($m=1$), the purple solid lines are the retrograde modes (m=1m=-1), and the orange dotted-dashed lines are the zonal modes. These are evaluated in the inertial frame of reference and so correspond to the frequencies observed through the brightness variations by a distant observer. The gray shaded region indicates the two months of strong accretion during the DNe outburst. The minus values of the time axis represent the time before the strong accretion event.
Both the panels exhibit relatively similar trends, as each radial mode order frequency gradually relaxes to its preoutburst value within approximately 36 months. The sudden temperature rise in the outer layers due to the accretion event causes the mode frequencies to rise by a few percent relative to their quiescent value. However, the lowest order $g$-modes (for example, $n_{\rm g}=2$) are somewhat less affected by the temperature change in the accreted envelope as they reside much deeper into the star, making their mode frequencies less sensitive to heating in the outer layers. While the overall qualitative behavior of mode relaxation remains consistent and similar with our earlier work (Kumar & Townsley, 2023), the WD models presented in this work incorporate significant improvements. The presence of element diffusion during the long-term accretion phase leads to smoother buoyancy profiles between the accreted material and the WD core. Additionally, our treatment of CV evolution is relatively realistic with adjusted core temperatures to reflect the observational constraints, hence giving a more correct mode frequency within the star.
@@
Figure 8: The rate of period change as a function of average observed periods of mode orders $n_{\rm g}=1-12$, following the strong intermittent accretion aka Dwarf Nova cycle for the several core temperatures (as labeled in the top panel) of a 0.78 $M_{\odot}$ model with an accreted layer mass of $1.5\times 10^{-4}~M_{\odot}$. The rates are evaluated at 4.5 months after the outburst. The dipole ($\ell=1$) modes, retrograde, zonal, and prograde modes are displayed in the top, middle, and bottom panels, respectively. Closed symbols indicate for the modes that are driven in the star’s frame and therefore likely to be observed. The minimum driving mode period is taken as 192 seconds, in agreement with our earlier work (Kumar & Townsley, 2025).
@@
Identifying the true mode orders in accreting WDs is challenging due to the limited number of observed modes. A common approach is to infer the radial order mode by identifying the consecutive radial order modes and matching them with the theoretical models within the reasonable parameter space. Here, we propose a potential new method for mode identification based on the time-variation of period or frequency during the months of cooling after the accretion event. Figure 8 presents the period derivative ($|\dot{P}|$) as a function of average observed periods for mode orders $n_{\rm g}=1-12$ for the dipole ($\ell=1$) modes, following the strong intermittent accretion event for $T_{\rm c}=5,6,6.5$, and $7\times 10^{6}$ K of a 0.78 $M_{\odot}$ model with $M_{\rm acc}=1.5\times 10^{-4}~M_{\odot}$. The period derivative is evaluated at 4.5 months post DN outbursts. A fiducial shortest driven mode period of approximately 192 seconds is adopted in star’s frame, a representative value that might be appropriate for GW Lib. Filled symbols indicate modes that are likely to be observed. Notably, retrograde modes exhibit the largest $|\dot{P}|$, approximately an order of magnitude higher than prograde and zonal modes. By observing a specific mode and computing its period derivative, this could reveal the true mode order without examining the neighboring modes. This technique may offer a valuable tool for mode identification.
@@
Figure 9 investigates the observed mode frequency variation of the first 14 radial orders (top to bottom) with the accreted layer mass, $M_{\rm acc}$, for a 0.78 $M_{\odot}$ WD model with $T_{\rm c}=5\times 10^{6}$ K. In general, mode frequencies increase with $M_{\rm acc}$. The solid purple, dashed blue, and dotted orange lines represent retrograde, prograde, and zonal modes for the $\ell=1$ triplets, respectively. The overall mode frequency increases with the newly accreted layer. The actual reason for this change is shown in Figure 10 - the Brunt, and therefore the WKB integrand, is higher in the envelope, so that as $M_{\rm acc}$ increases under the integral curve shown in Figure 10 increases, and the mode frequencies are proportional to this integral, as discussed in the previous subsection. While the $n_{\rm g}=1$ mode exhibits the largest change and continuous increase with $M_{\rm acc}$, the other radial order modes show a slight dip when $M_{\rm acc}$ approaches the $M_{\rm ign}$. The evolving accreted envelope effectively breaks the star into three distinct cavities: core, shell, and transition layer. The transition layer between the core and the shell plays a crucial role in determining node locations for different mode orders. While the mode orders $n_{\rm g}>1$ warrant the existence of at least one node either in the shell or core, the single node in the $n_{\rm g}=1$ mode constrained near the transition layer, making it highly sensitive to variations in the newly accreted layer.
@@
A striking feature that is quite evident in Figure 9 is mode pairing (alternating small-large frequency spacing). This is particularly noticeable in the both prograde and retrograde mode orders, such as prograde $n_{\rm g}=3-4$, 565-6 and retrograde $n_{\rm g}=6-7$. As the radial order increases, a node is allocated either in the core or the envelope, depending on the mode. This leads to alternation between the smaller and larger frequency spacings between the consecutive modes. In a uniform stellar structure, $g$-modes are expected to have an inverse spacing in frequency, with higher-order modes appearing closer together. However, in WDs, the contrast between the light-element envelope and the heavier-element core disrupts this regularity making the spacing irregular (Townsley et al., 2004; Fontaine & Brassard, 2008; Althaus & Córsico, 2022). In our accreting WD, there is a single localized division between the accreted light-element (H+He) layer and the inner CO material, since we have hypothesized that the entire envelope will be ejected and then re-accreted. The thickness and structure of the subsurface layer modify the density gradient, therefore shifting the Brunt-Väisälä frequency profile. These changes are significantly effective in the transition region between the core and envelope, where the mode confinement and trapping conditions are set. The modification to Buoyancy frequency appears directly in the WKB integrand, which determines the strength of the cavities.
@@
Having discussed avoided mode crossings, we can now remark more fully upon a feature seen as $T_{\rm c}$ was varied in section 3.1. This features differs from a traditional “avoided” mode crossing, resembling more of a “super” mode crossing. As $T_{\rm c}$ increases in Figure 6 one of the nodes in the $n_{\rm g}=5$ mode moves from the boundary region into the envelope, and one of the nodes in the $n_{\rm g}=3$ mode move from the core into the boundary region. This is likely so strong partially because of the change in boundary condition due to the presence and then absence of the solid core.
Figure 11 demonstrates the period derivative ($|\dot{P}|$), calculated between three and six months following a dwarf nova outburst, as a function of average observed mode periods for the dipole triplets of radial orders $n_{\rm g}=1-15$. The WD models shown correspond to accreted layer masses of $M=1.63,2.06,2.48,2.81,3.05$, and $3.47\times 10^{-4}M_{\odot}$ for a 0.78 $M_{\odot}$ with a fixed $T_{\rm c}=5~\times~10^{6}$ K. All models have the same final $T_{\rm eff}=14000$ K. With respect to a shortest driven period of 192 s in the star’s frame, the filled and open symbols indicate whether the modes are likely to be observed or not. The retrograde modes reveal larger $|\dot{P}|$ values compared to their prograde and zonal counterparts. While the lower-order $g$ modes show minimum sensitivity in $|\dot{P}|$ to changes in $M_{\rm acc}$, the higher-order modes display discernible variations. Particularly, the sudden decrease in $|\dot{P}|$ for the prograde mode orders $n_{\rm g}=5-6$ can be attributed to the previously discussed “avoided mode crossing” behavior, as illustrated in Figure 9. A similar feature is also observed for the retrograde mode orders $n_{\rm g}=6-7$, as seen in the top panel of Figure 11. That is, one of these modes is the “trapped” envelope mode, and will thus have a high $\dot{P}$, and other is not. Within the framework of the known mode driving mechanism for accreting WDs (Arras et al., 2006), and considering the shortest driven period adopted in this work, we find that nearly all retrograde $g$ modes up to radial order 12 are expected to be driven and thus observable, whereas only the lowest order $n_{\rm g}=1$ prograde mode is likely to be excited, as illustrated in Figure 11. This suggests that retrograde $g$ modes are the most likely candidates for detection in accreting WDs and similar systems (Kumar & Townsley, 2025).
@@
Figure 13 explores $|\dot{P}|$ as a function of observed mode period for various WD mass models following a dwarf-novae event, similar to Fig. 11. Lower WD models possess longer retrograde mode periods and also show greater $|\dot{P}|$. In particular, for the 0.60 $M_{\odot}$ model, mode orders $n_{\rm g}=4$ to 7 show a continuous increase to $|\dot{P}|$, peaking at $n_{\rm g}=7$ in both retrograde and prograde modes. This behavior, however, is not evident in the higher mass models. The lower-order modes show the least variations to $|\dot{P}|$ across all the WD models.
@@ 4 Discussion and Conclusions
The seismology of WDs in CVs is relatively less well understood than that of isolated WDs due to several challenges, including their being subject to frequent outbursts, very faint observational signatures, and the intrinsic difficulty in modeling their exact evolution (Schreiber et al., 2016; Shen & Quataert, 2022). In this work, we have presented the first comprehensive seismological analysis of the accreting WDs using $g$ modes. We use a forward modeling approach rather than trying to fit the observed modes in a particular star. To this end, we choose representative values of the parameters characterizing the WD. Specifically, we investigate three main modeling components that shape these systems: the core temperature ($T_{\rm c}$), the mass of the accreted H-rich layer ($M_{\rm acc}$), and the WD mass ($M_{\rm WD}$). We envisage that $g$ modes are useful for probing these properties in rapidly rotating accreting WDs. While previous studies proposed that Rossby modes ($r$ modes) may be more favorable in such systems (Saio, 2019), recent work (Kumar & Townsley, 2025) demonstrated that $g$ modes are more likely to dominate pulsations in accreting WDs.
@@
We compute $g$ mode frequencies for low-order modes during quiescence as well as their variation following a dwarf nova outburst for the underlying WD models summarized in Table 1. The period spacings change in distinctive ways for each variation in the parameter of the WD. The variation of $T_{\rm c}$ seems to show the most change in which order mode shows minimum period spacing. The variations of $M_{\rm acc}$ have the largest effect on the $n_{g}=1$ mode and lead to modest changes in the order of the mode with minimum mode spacing. A change in mass leads mostly to a general increase in all mode frequencies, but low-order modes are the most affected. Furthermore, the presence or absence of a solid core doesn’t seem to have a strong effect, based either on variation of $T_{\rm c}$ or mass, but it shows some potential significance. The higher-order modes show more effect compared to the lower ones with the extent of crystallization, as they do not follow the asymptotic prescriptions, $\Pi\propto(\int_{r1}^{r2}|N|dr/r)^{-1}$ (Tassoul et al., 1990). This is explained by using the WKB integrand shown in Figure 7, as the envelope to core integrand reduces from lower to higher $T_{\rm c}$. The fact that some radial orders frequencies are strongly affected by the changes in the extent of the crystallized core while others are not, indicates that we may be able to possibly unwind the impact of crystallization with the help of various surface layer masses.
@@
The general and qualitative trend of a particular mode after the dwarf nova remains similar to what is presented in (Kumar & Townsley, 2023), relaxing to its pre-outburst value in a few month’s time. However, the WD models presented here have significant improvements from our previous work.

### v2

@@ 2.5 Choice of Modes
The associated pulsations in accreting WDs are generally attributed to be $g$ modes similar to those commonly applied to the isolated WDs. These $g$ modes are excited thermally by the convection zone related to hydrogen and helium ionization zones, as in DAV and DBV stars (Brickhill, 1991; Arras et al., 2006; Saio, 2013). However, some recent studies in the literature suggest that these pulsations could be Rossby modes (Saio, 2019). In our companion paper, we conducted a simultaneous analysis of both $g$ and $r$ modes, demonstrating that within the currently established framework of the convective driving mechanism, the pulsations in accreting WDs should be expected to have a $g$-mode origin (Kumar & Townsley, 2026). We explicitly computed the visibility amplitudes of both $g$ and $r$ modes and highlighted that while a few retrograde $r$ modes exhibit larger visibility, the low-order $g$ modes possess higher frequencies in the star’s frame, thus making them more likely to be driven in the star’s frame and observable.
Following the findings of Kumar & Townsley (2026), with the investigation of the surface amplitudes of $g$ modes in accreting WDs, we focus exclusively on $g$ modes up to the second-lowest order ($\ell=2$) and disregard the $r$ mode significance in such systems in this present analysis. We use GYRE to compute adiabatic $g$-mode frequencies just as described in Kumar & Townsley (2026).
@@ 3 Results
We now present our seismological calculations for the accreting WD models described in section 2. We consider the $g$ modes in the rapid rotation limit using the observed spin period of GW Lib, 209 s, as a representative value for CV systems (Szkody et al., 2012). This period is much shorter than the hours or days observed in isolated WDs. In this regime, rotation starts to strongly and nonlinearly modify $g$-mode frequencies once the spin parameter $2\Omega/\omega\gtrsim 1$. Given that $g$-mode periods in GW Lib are comparable to its spin period, the system lies within the regime where rotational effects are significant and hence the Traditional Approximation of Rotation (TAR) is needed. Although, GYRE allows the treatment of rotation both with and without the Cowling approximation (Cowling, 1941), we conduct analysis within the TAR and adopt the Cowling approximation, consistent with common practice in the literature (Bildsten et al., 1996; Lee & Saio, 1987; Townsend, 2020). By disabling it in GYRE, we find that $g$-mode frequencies will change by about 0.002% if the Cowling approximation is not used. We present $g$-mode frequency variation with $T_{\rm c}$ in section 3.1, quantifying the mode frequencies with and without a partial solid core. Furthermore, we propose a potential new method to identify the mode orders based on the time variation of frequencies during the months of cooling after the strong accretion event. Section 3.2 introduces the variation of the $g$-mode frequencies with the accreted layer of the star, $M_{\rm acc}$. Finally, we show the changes for the different mass models.
@@
Here we begin by showing the $g$-mode eigenfrequencies following the strong accretion event. These strong accretion outbursts occur on relatively short timescales compared to the overall age of the star. However, the intermittent accretion events with finite quiescence periods can significantly disrupt the outer layers, leading to modifications in the $g$-mode frequencies. We investigate the radial mode orders for the dipole mode ($\ell=1$) after the strong accretion event, tracking the evolution of mode frequencies over a period of just over four years. These calculations build upon the work of Kumar & Townsley (2023), but with a significantly improved WD model as outlined in the section 2.2. Figure 4 illustrates the time variations of the $g$-mode frequencies for radial orders $n_{\rm g}=2-16~(top~to~bottom)$ in the observer’s frame of a dipole mode, following the accretion outbursts for a 0.78 $M_{\odot}$ WD model. The left panel corresponds to $T_{\rm c}=5\times 10^{6}$ K, where the WD core has a solid core, while the right panel shows results for $T_{\rm c}=7\times 10^{6}$ K. The dashed blue lines are the prograde modes ($m=1$), the purple solid lines are the retrograde modes (m=1m=-1), and the orange dotted-dashed lines are the zonal modes. These are evaluated in the inertial frame of reference and so correspond to the frequencies observed through the brightness variations by a distant observer. The gray shaded region indicates the two months of strong accretion during the DNe outburst. The minus values of the time axis represent the time before the strong accretion event.
Both the panels exhibit relatively similar trends, as each radial mode order frequency gradually relaxes to its preoutburst value within approximately 36 months. The sudden temperature rise in the outer layers due to the accretion event causes the mode frequencies to rise by a few percent relative to their quiescent value. However, the lowest order $g$ modes (for example, $n_{\rm g}=2$) are somewhat less affected by the temperature change in the accreted envelope as they reside much deeper into the star, making their mode frequencies less sensitive to heating in the outer layers. While the overall qualitative behavior of mode relaxation remains consistent and similar with our earlier work (Kumar & Townsley, 2023), the WD models presented in this work incorporate significant improvements. The presence of element diffusion during the long-term accretion phase leads to smoother buoyancy profiles between the accreted material and the WD core. Additionally, our treatment of CV evolution is relatively realistic with adjusted core temperatures to reflect the observational constraints, hence giving a more correct mode frequency within the star.
@@
Figure 8: The rate of period change as a function of average observed periods of mode orders $n_{\rm g}=1-12$, following the strong intermittent accretion aka Dwarf Nova cycle for the several core temperatures (as labeled in the top panel) of a 0.78 $M_{\odot}$ model with an accreted layer mass of $1.5\times 10^{-4}~M_{\odot}$. The rates are evaluated at 4.5 months after the outburst. The dipole ($\ell=1$) modes, retrograde, zonal, and prograde modes are displayed in the top, middle, and bottom panels, respectively. Closed symbols indicate for the modes that are driven in the star’s frame and therefore likely to be observed. The minimum driving mode period is taken as 192 seconds, in agreement with our earlier work (Kumar & Townsley, 2026).
@@
Identifying the true mode orders in accreting WDs is challenging due to the limited number of observed modes. A common approach is to infer the radial order mode by identifying the consecutive radial order modes and matching them with the theoretical models within the reasonable parameter space. Here, we propose a potential new method for mode identification based on the time-variation of period or frequency during the months of cooling after the accretion event. Figure 8 presents the period derivative ($|\dot{P}|$) as a function of average observed periods for mode orders $n_{\rm g}=1-12$ for the dipole ($\ell=1$) modes, following the strong intermittent accretion event for $T_{\rm c}=5,6,6.5$, and $7\times 10^{6}$ K of a 0.78 $M_{\odot}$ model with $M_{\rm acc}=1.5\times 10^{-4}~M_{\odot}$. The period derivative is evaluated at 4.5 months post DN outbursts. A fiducial shortest driven mode period of approximately 192 seconds is adopted in star’s frame, rather than deriving it from a full non-adiabatic mode calculation; this value is representative of what may be appropriate for GW Lib. Filled symbols indicate modes that are likely to be observed. Notably, retrograde modes exhibit the largest $|\dot{P}|$, approximately an order of magnitude higher than prograde and zonal modes. By observing a specific mode and computing its period derivative, this could reveal the true mode order without examining the neighboring modes. This technique may offer a valuable tool for mode identification.
@@
Figure 9 investigates the observed mode frequency variation of the first 14 radial orders (top to bottom) with the accreted layer mass, $M_{\rm acc}$, for a 0.78 $M_{\odot}$ WD model with $T_{\rm c}=5\times 10^{6}$ K. In general, mode frequencies increase with $M_{\rm acc}$. The solid purple, dashed blue, and dotted orange lines represent retrograde, prograde, and zonal modes for the $\ell=1$ triplets, respectively. The overall mode frequency increases with the newly accreted layer. The actual reason for this change is shown in Figure 10 - the Brunt, and therefore the WKB integrand, is higher in the envelope, so that as $M_{\rm acc}$ increases, the area under the integral curve shown in Figure 10 increases. This leads to higher mode frequencies since the mode frequencies are proportional to this integral, as discussed in the previous subsection. This enhancement arises both from hydrogen to helium transition in the Brunt-Väisälä frequency and from the structural changes to the outer cavity associated with the newly accreted layer. While the $n_{\rm g}=1$ mode exhibits the largest change and continuous increase with $M_{\rm acc}$, the other radial order modes show a slight dip when $M_{\rm acc}$ approaches the $M_{\rm ign}$. The evolving accreted envelope effectively breaks the star into three distinct cavities: core, shell, and transition layer. The transition layer between the core and the shell plays a crucial role in determining node locations for different mode orders. While the mode orders $n_{\rm g}>1$ warrant the existence of at least one node either in the shell or core, the single node in the $n_{\rm g}=1$ mode constrained near the transition layer, making it highly sensitive to variations in the newly accreted layer.
@@
A striking feature that is quite evident in Figure 9 is mode pairing (alternating small-large frequency spacing). This is particularly noticeable in the both prograde and retrograde mode orders, such as prograde $n_{\rm g}=3-4$, 565-6 and retrograde $n_{\rm g}=6-7$. As the radial order increases, a node is allocated either in the core or the envelope, depending on the mode. This leads to alternation between the smaller and larger frequency spacings between the consecutive modes. In a uniform stellar structure, $g$ modes are expected to have inverse frequencies which are uniformly spaced, with higher-order modes appearing closer together. However, in WDs, the contrast between the light-element envelope and the heavier-element core disrupts this regularity making the spacing irregular (Townsley et al., 2004; Fontaine & Brassard, 2008; Althaus & Córsico, 2022). In our accreting WD, there is a single localized division between the accreted light-element (H+He) layer and the inner CO material, since we have hypothesized that the entire envelope will be ejected and then re-accreted. The thickness and structure of the subsurface layer modify the density gradient, therefore shifting the Brunt-Väisälä frequency profile. These changes are significantly effective in the transition region between the core and envelope, where the mode confinement and trapping conditions are set. The modification to Buoyancy frequency appears directly in the WKB integrand, which determines the strength of the cavities.
@@
Having discussed avoided mode crossings, we can now remark more fully upon a feature seen as $T_{\rm c}$ was varied in section 3.1. This feature differs from a traditional “avoided” mode crossing, resembling more of a “super” mode crossing. As $T_{\rm c}$ increases in Figure 6 one of the nodes in the $n_{\rm g}=5$ mode moves from the boundary region into the envelope, and one of the nodes in the $n_{\rm g}=3$ mode move from the core into the boundary region. This is likely so strong partially because of the change in boundary condition due to the presence and then absence of the solid core.
Figure 11 demonstrates the period derivative ($|\dot{P}|$), calculated between three and six months following a dwarf nova outburst, as a function of average observed mode periods for the dipole triplets of radial orders $n_{\rm g}=1-15$. The WD models shown correspond to accreted layer masses of $M=1.63,2.06,2.48,2.81,3.05$, and $3.47\times 10^{-4}M_{\odot}$ for a 0.78 $M_{\odot}$ with a fixed $T_{\rm c}=5~\times~10^{6}$ K. All models have the same final $T_{\rm eff}=14000$ K. With respect to a shortest driven period of 192 s in the star’s frame, the filled and open symbols indicate whether the modes are likely to be observed or not. The retrograde modes reveal larger $|\dot{P}|$ values compared to their prograde and zonal counterparts. While the lower-order $g$ modes show minimum sensitivity in $|\dot{P}|$ to changes in $M_{\rm acc}$, the higher-order modes display discernible variations. Particularly, the sudden decrease in $|\dot{P}|$ for the prograde mode orders $n_{\rm g}=5-6$ can be attributed to the previously discussed “avoided mode crossing” behavior, as illustrated in Figure 9. A similar feature is also observed for the retrograde mode orders $n_{\rm g}=6-7$, as seen in the top panel of Figure 11. That is, one of these modes is the “trapped” envelope mode, and will thus have a high $\dot{P}$, and other is not. Within the framework of the known mode driving mechanism for accreting WDs (Arras et al., 2006), and considering the shortest driven period adopted in this work, we find that nearly all retrograde $g$ modes up to radial order 12 are expected to be driven and thus observable, whereas only the lowest order $n_{\rm g}=1$ prograde mode is likely to be excited, as illustrated in Figure 11. This suggests that retrograde $g$ modes are the most likely candidates for detection in accreting WDs and similar systems (Kumar & Townsley, 2026).
@@
Figure 13 explores $|\dot{P}|$ as a function of observed mode period for various WD mass models following a dwarf-novae event, similar to Fig. 11. Lower-mass models possess longer retrograde mode periods and also show greater $|\dot{P}|$. In particular, for the 0.60 $M_{\odot}$ model, mode orders $n_{\rm g}=4$ to 7 show a continuous increase to $|\dot{P}|$, peaking at $n_{\rm g}=7$ in both retrograde and prograde modes. This behavior, however, is not evident in the higher mass models. The lower-order modes show the smallest variations to $|\dot{P}|$ across all the WD models.
@@ 4 Discussion and Conclusions
The seismology of WDs in CVs is relatively less well understood than that of isolated WDs due to several challenges, including their being subject to frequent outbursts, very faint observational signatures, and the intrinsic difficulty in modeling their exact evolution (Schreiber et al., 2016; Shen & Quataert, 2022). In this work, we have presented the first comprehensive seismological analysis of the accreting WDs using $g$ modes. We use a forward modeling approach rather than trying to fit the observed modes in a particular star. To this end, we choose representative values of the parameters characterizing the WD. Specifically, we investigate three main modeling components that shape these systems: the core temperature, the mass of the accreted H-rich layer ($M_{\rm acc}$), and the WD mass. We envisage that $g$ modes are useful for probing these properties in rapidly rotating accreting WDs. While previous studies proposed that Rossby modes ($r$ modes) may be more favorable in such systems (Saio, 2019), recent work (Kumar & Townsley, 2026) demonstrated that $g$ modes are more likely to dominate pulsations in accreting WDs.
@@
We compute $g$-mode frequencies for low-order modes during quiescence as well as their variation following a dwarf nova outburst for the underlying WD models summarized in Table 1. The period spacings change in distinctive ways for each variation in the parameter of the WD. The variation of $T_{\rm c}$ seems to show the most change in which order mode shows minimum period spacing. The variations of $M_{\rm acc}$ have the largest effect on the $n_{g}=1$ mode and lead to modest changes in the order of the mode with minimum mode spacing. A change in mass leads mostly to a general increase in all mode frequencies, but low-order modes are the most affected. Furthermore, the presence or absence of a solid core doesn’t seem to have a strong effect, based either on variation of $T_{\rm c}$ or mass, but it shows some potential significance. While the higher-order $g$ modes generally follow the asymptotic relation, $\Pi\propto(\int_{r1}^{r2}|N|dr/r)^{-1}$ (Tassoul et al., 1990), deviations from this relation are still apparent. This indicates that the crystallization and structural changes modify the finite departure from the ideal asymptotic relation. This is explained by using the WKB integrand shown in Figure 7, as the envelope to core integrand decreases as $T_{\rm c}$ increases, leading to systematic changes in $N/r$. The fact that some radial orders frequencies are strongly affected by the changes in the extent of the crystallized core while others are not, indicates that we may be able to possibly unwind the impact of crystallization with the help of various surface layer masses.
We note that more recent studies have proposed that the WD crystallization may generate magnetic fields (Isern et al., 2017), potentially transforming CV to magnetic systems (Schreiber et al., 2021). While current models suggest that thermohaline mixing associated with crystallization is too weak to sustain a strong magnetic dynamo (Montgomery & Dunlap, 2024), convective motions may still be able to dredge-up previously buried magnetic fields (Blatman & Ginzburg, 2024a, b) and the presence of such magnetic fields could, in turn, suppress the lower-frequency $g$ modes (Rui et al., 2025). However, in this work, we do not consider any magnetic effects, and our results therefore apply most directly to non-magnetic CVs. A self-consistent treatment of crystallization-driven magnetism and its impact on pulsation modes should be investigated in detail for accreting systems in future work.
@@
The general and qualitative trend of a particular mode after the dwarf nova remains similar to what is presented in Kumar & Townsley (2023), relaxing to its pre-outburst value in a few months time. However, the WD models presented here have significant improvements from our previous work.
@@
We also note that the low-frequency $g$ modes in WDs can exhibit stochastic variations in their frequencies (Montgomery et al., 2010; Hermes et al., 2017), and this variability may introduce additional scatter in the observed mode periods. As mentioned by Hermes et al. (2017) for the DAV case, such variability is mostly seen with periods larger than 800 second. Those observations pertain to pure hydrogen atmosphere WDs, whereas the system of interest here involve accreting white dwarfs with mixed atmosphere, which might further complicate things. Despite this, we believe that the overall trends underlying our proposed observables should remain largely unchanged.
The relevant input files and datasets are available online at doi: https://doi.org/10.5281/zenodo.18307075